OpenStudy (anonymous):

An object is projected upward at a rate of 160 feet per second. Find the time when it is 384 feet above the ground.. The equation given is h=-16t^2+V(o)t. 384=-16*t^2+160*t

OpenStudy (btaylor):

Good. Now put it all one one side of the equation, and use the quadratic formula. You should get t=4 and 6

OpenStudy (anonymous):

-16*t^2+160*t-384=0?

OpenStudy (btaylor):

Yep. Now, you can divide by -16 to get t^2-10t+24=0. This factors into (t-6)(t-4)=0, so you get t=6 and 4

OpenStudy (anonymous):

Thanks!